Prove that if a sequence {an}∞ n=1 satisfies Cauchy’s criterion, then it is bounded.


1
Expert's answer
2021-02-24T12:37:47-0500

Suppose "(a_n)" is a sequence satisfied Cauchy's criterion .

Claim : "(a_n)" is bounded i,e, there exit a "C\in \N" such that "|a_n|\leq C" for all "n\in \N" .

Since "(a_n)" is a Cauchy's Sequence , therefore for any "\epsilon >0" there exist a "K\in \N" such that "| a_n-a_m|<\epsilon" "\forall m,n > K" .


We know that ,

"|a_n| = | a_n+a_m-a_m| \leq |a_m|+|a_n-a_m|" by triangular inequality .

Set "\epsilon =1" , Because this is cauchy ,

There exist "K\in \N" such that "|a_n-a_m|\leq 1" "\forall m,n >K"

Set "m=K+1" , Combined with our initial note , we can re write the following ,

"|a_n|\leq 1+|a_{K+1}|" and this is true for all "n>K"

This bound all the term beyond the "Kth" .

Looking at the term before the "Kth" term , we can find the maximum of them and notes that this bounds that part of the sequence

"|a_n| \leq Max \{ |a_1|,|a_2|,.........,|a_K|\}"

And this is true for all "n\leq K" .

By choosing maximum of either "1+|a_{K+1}|" or the maximum of the aforementioned set we can find our "C"

Which bound all the term of the sequence .